# Model for Peanuts: Hijacking ML models without Training Access is Possible

### v1

@@ 10.1. Meta-unlearning: meta-Learning-based malicious task unlearning
The intuition is that some internal representations are more transferable across tasks than others. For example, given an original task $\mathcal{T}_{i}$ and a related task $\mathcal{T}_{j}$, i.e., both $\mathcal{T}_{i},{T}_{j}\sim p(\mathcal{T})$, where $p(\mathcal{T})$ is distribution of tasks, a model might learn internal features that are broadly applicable to all tasks in $p(\mathcal{T})$, and others that are specific to the original task $\mathcal{T}_{i}$. Our objective is to penalize the learning of hijacking task-specific features while learning the ones relevant to the original one. The proposed approach is inspired by the meta-learning literature (Finn et al., 2017): We train the model on $\mathcal{T}_{i}$ while maximizing the loss function $\mathcal{L}_{j}$, relative to the (hijacking) task $\mathcal{T}_{j}$. The approach is detailed in Algorithm 1. Tables 2 and 3 show the results of experiments designed to test the effectiveness of the meta-unlearning to retain accuracy on the original classification task while resisting SnatchML. We also add a hijacking setting with stronger access than our approach, assuming the hijacking dataset can be used to train a neural network for hijacking (NN Surrogate), not only using distances such as in SnatchML. Table 3 shows the results for gender classification as the original task and ethnicity recognition as the hijacking task. It shows that when the model is trained with unlearning, the accuracy for the original task generally decreases. This is likely a consequence of the unlearning procedure, which may also remove features beneficial to the original task. While the surrogate NN achieves, as expected, higher hijacking success, we notice that the meta-unlearning is more efficient on this attack than SnatchML, suggesting a higher robustness of distance-based inference in this setting. For example, the defense has not significantly impacted the accuracy of the hijacking task (Ethnicity). After empirically exploring the loss weights, these results were obtained with $\alpha=1$ and $\beta=0.01$. Increasing the unlearning loss coefficient leads to an unacceptable accuracy drop in the original task.
Table 2. Accuracy of the original task (ER) and hijacking attack (re-identification) with and without meta-unlearning.
@@
Table 3. Accuracy of the original task (Gender) and hijacking attack (Ethnicity) with and without meta-unlreaning.
@@
Table 4. Accuracy of the original task (PDD) and hijacking attack (Pulmonary infection type recognition) with and without meta-unlearning.
@@
Table 5. Model compression: Results report the accuracy of the original and hijacking tasks on different datasets and models. For each case, we report accuracy values for the baseline model $f(.)$ and compact model $f_{cmp}(.)$ found by our exhaustive search for width expansion ratios ($m$). We evaluate the hijacking under black-box (Logits) and white-box (FV) settings.
[…]
Resnet-9 0.937 0.935 0.1$\times$ 0.257 0.217 0.293 0.294
Mobilenet 0.939 0.915 0.45$\times$ 0.148 0.143 0.176 0.163
Resnet-9 0.937 0.935 0.1$\times$ 0.232 0.144 0.479 0.359
Mobilenet 0.939 0.915 0.45$\times$ 0.074 0.070 0.076 0.066
Resnet-9 0.853 0.851 0.1$\times$ 0.764 0.541 0.787 0.756
Mobilenet 0.801 0.780 0.1$\times$ 0.544 0.482 0.751 0.633
10.2. Model Compression as a Defense
The previous approach is limited by the constraint of known hijacking tasks, which may not always reflect real-world scenarios. We aim to overcome this limitation in this second approach by proposing a more task-agnostic methodology. In Section 9.1, we observed a correlation between the model size and the effectiveness of SnatchML. In this section, we explore the extent to which model compression can help defend against SnatchML.
(4) $f(.)=l^{n}\circ l^{n-1}\circ l^{n-2}\circ\dots\circ l^{2}\circ l^{1}s.t.\;\;f\in\mathcal{F}$
(5) Where for each $\text{Where for each layer}\quad l^{j}=\{W_{1}^{j},W_{2}^{j},W_{3}^{j},...,W_{m_{j}}^{j}\}$
(6) $f_{cmp}=\operatorname*{arg\,min}_{f_{m}\in\mathcal{F}}\alpha\cdot\mathcal{L}_{i}(f_{m}(X),Y)+\beta\cdot\mathcal{C}ount(f_{m})$
Experimental Setup. We design a search space of $14$ candidate width expansion ratios, ranging from 0.1×0.1$\times$0.1 × to 0.75×0.75$\times$0.75 ×. For each model $f$, we exhaustively iterate through the search space and sample an expansion ratio $m$ to get a compact model $f_{m}(.)$. At each iteration, we evaluate $f_{m}(.)$ using the loss $\mathcal{L}_{i}(.)$ (as a proxy for utility on the original task) and parameters count $\mathcal{C}ount(.)$ (as a proxy for model size). After the exhaustive search, we rank all candidate $f_{m}(.)$ according to their loss values and parameters count using the TOPSIS method (Behzadian et al., 2012) to render an optimal $f_{cmp}(.)$ with a tradeoff between utility and model size.
Results and Discussion. To test if lowering the models’ dimensionality enhances their resistance to hijacking, we evaluate our compression method on three hijacking attacks: user re-identification on CK (Lucey et al., 2010) identification on Olivetti (Samaria and Harter, 1994), and Pneumonia recognition on Chest X-ray (Kermany et al., 2018). Results are detailed in Table 5 for each hijacking attack using three models: 2D-CNN, ResNet-9, and MobileNet. Across the three datasets, compact models showed a notable reduction in hijacking task accuracy. Specifically, ResNet-9 compact models demonstrate high resilience to hijacking while maintaining comparable accuracy to the baseline model. For example, identification accuracy on Olivetti and pneumonia type recognition on Chest X-ray decreases by $29\%$ and $38\%$, respectively, even under an aggressive compression strategy with a width expansion of $m=0.1\times$ on the baseline. Conversely, MobileNet shows a modest reduction in hijacking attack vulnerability for re-identification and identification tasks at the expense of a slight decrease in utility for the original task of ER (approximately $\sim$3reduction). This finding is consistent with the discussions in Section 9.1, highlighting the intrinsic compactness of the baseline MobileNet model. Notably, in the pneumonia scenario, MobileNet with a $m=0.1\times$ shows low hijacking accuracy with a decrease of $11.3\%$ (Logits) and $15.7\%$ (FV).
It’s worth noting that model compression does not always reduce the hijacking accuracy in specific scenarios, e.g., re-identification on CK+. This phenomenon may be attributed to the overlap between the original and hijacking tasks, particularly in smaller datasets where models are prone to overfitting. Such models tend to capture fine-grained features necessary for the original task, which concurrently exhibit a high correlation with the hijacking task. Isolating these fine-grained features could compromise the utility of the original task. Therefore, more sophisticated techniques of learning combined with model compression are needed.
11. Discussion and Concluding remarks
Increased potential for harm with less access privilege. In this paper, we propose a new threat model within the model hijacking attack scenario. Specifically, we establish a new risk of adversaries hijacking ML models with restricted access, i.e., without access to the training process or data. We propose SnatchML, which uses features extracted by the victim model at inference time to infer a different task defined by the attacker. In contrast to related hijacking attacks, where those of the original task limit the number of classes in the hijacking task, our proposed approach imposes no predefined constraints on the number of classes involved in the hijacking task. Therefore, SnatchML  not only assumes a stronger threat model but also enables attackers with lower access privileges to potentially cause greater harm.
Insights towards security-aware NAS. We suggest that the vulnerability of ML models to this type of attack is primarily due to their capacity to unintentionally learn ”extra-knowledge”, which is mainly related to over-parametrization. Interestingly, our results reveal a disparity in attack success based on the model architecture. This observation could be particularly significant in the Neural Architecture Search towards developing secure-by-design models.
Beyond the technical implications of SnatchML. We believe SnatchML  challenges some of the AI regulatory efforts’ foundations and underscores the need for more comprehensive approaches. The potential for compliant ML models to be repurposed for unethical or illegal tasks poses significant accountability risks for model owners. This highlights a critical oversight in the EU AI Act and similar regulatory frameworks, which assume that models trained on specific tasks are confined to those boundaries.
@@ 11. Discussion and Concluding remarks
Attack Limitations. In our study, we demonstrate that the hijacking effectiveness of SnatchMLcan reach state-of-the-art task accuracy in some cases. However, this effectiveness can be influenced by several factors, particularly the victim model’s architecture. This limitation does not exist in traditional hijacking attacks where the attacker typically has access to training data, allowing them to directly manipulate the model’s parameters. Despite this limitation, our approach represents a significant advancement over traditional hijacking methods by overcoming the constraints related to the number of classes in the hijacking task.
Defenses Limitations. The meta-unlearning defense strategy we propose is designed to unlearn or forget specific tasks that are known to pose ethical or privacy concerns. However, one significant limitation of this approach is its specificity to the task it is designed to counter. This makes it less useful in environments where the potential hijacking tasks are unknown or can vary widely. However, we believe more advanced meta-learning-inspired approaches that define a distribution of tasks instead of one could be more efficient.
The compression defense mechanism we explore also presents its challenges, particularly when dealing with tasks with semantic or functional overlap with the original task. The necessity for aggressive model compression to disrupt the latent space, where features might otherwise be exploited for a hijacking task, can inadvertently reduce the model’s accuracy on its original task. While effectively reducing the model’s vulnerability to hijacking, this approach requires a delicate balance between maintaining the original task’s utility and mitigating the risk of hijacking. Exploring additional compression techniques, such as quantization and pruning followed by fine-tuning on the original dataset, could offer further robustness to the models.
In summary, we believe this work contributes a new perspective on the ML security landscape and calls for advancements in security-aware model design and evaluation.

### v2

@@ XI-B Meta-unlearning:
TABLE III: Accuracy of the original task (ER) and hijacking attack (re-identification) with and without meta-unlreaning.
@@
TABLE IV: Accuracy of the original task (Gender) and hijacking attack (Ethnicity) with and without meta-unlreaning.
@@
TABLE V: Impact of Meta-Unlearning on Original Task and Hijacking Task Accuracy
@@
XII Discussion
Potential for harm with less access privilege. In this paper, we introduce a novel threat model within the context of model hijacking attacks. Specifically, we identify a new risk where adversaries with very restricted access can hijack ML models. Unlike existing hijacking attacks, which are constrained by the number of classes in the original task, our approach imposes no such limitations on the hijacking task’s class count. Consequently, SnatchML  not only assumes a more robust threat model but also enables attackers with lower access privileges to inflict greater harm. Moreover, SnatchML  does not require input modification during inference, facilitating a more covert form of hijacking. This allows a malicious model owner to repurpose the model while ostensibly complying with AI regulations, as the attack can be executed using a model from a trusted vendor without altering the input data.
Beyond the technical implications. We contend that our work has also significant implications for risk-based regulatory frameworks, as it challenges some of their foundational assumptions. In fact, the debate on regulating AI-powered systems has gained global momentum ultimately exemplified by initiatives like the European Commission’s proposal for a risk-assessment framework known as the EU AI-Act [10]. This framework seeks to categorize potential risks based on the nature of the task for which the ML model is trained, i.e., the original task. The reliance on the learned task’s criticality as a risk metric is inevitably tied to the following implicit hypothesis: ”If an ML model is trained on a data distribution $\mathcal{D}$, to learning a task $\mathcal{T}$, it is unlikely to (unintentionally) learn another task $\mathcal{T^{\prime}}$”. Our work shows that it is not sufficient for a benign model to be securely trained on a legitimate original task to guarantee that it will be immune to repurposing for unethical or illegitimate tasks. Therefore, additional safeguards and assurances are necessary.
@@ XII Discussion
Attack Limitations. In our study, we demonstrate that the hijacking effectiveness of SnatchML can reach state-of-the-art task accuracy in some cases. However, this effectiveness can be influenced by several factors, particularly the victim model’s architecture. Because of its zero-shot aspect, the performance of SnatchML can be significantly affected by the distribution of the hijacking task, (inter-class/intra-class variance). For instance, in scenarios like ethnicity classification, different ethnic groups might have significant intra-class variance. Despite this limitation, our approach represents a significant advancement over traditional hijacking methods by overcoming the constraints related to the number of classes in the hijacking task.
XIII Related work
Model hijacking. Model hijacking has emerged as a critical concern in the deployment of ML systems. Traditional hijacking techniques typically necessitate access to the model’s training process, allowing for the insertion of malicious alterations during training to facilitate the hijacking. Several papers proposed to overload ML models with secondary tasks. Salem et al. [38] proposed Model Hijacking attack that hides a model covertly while training a victim model. Mallya et al. [30] proposed Packnet that trains the model with multiple tasks. Si et al. [41], propose the Ditto attack, which hijacks text generation models by camouflaging hijacking datasets to resemble original ones, focusing on the output rather than input, making the hijack difficult to detect. The authors in [38] introduce the Chameleon and Adverse Chameleon attacks, which utilize an encoder-decoder model to camouflage hijacking datasets to appear like the original dataset, allowing stealthy hijacking of machine learning models.
The main difference with our work is that these attacks assume a threat model similar to poisoning attacks with the attacker having access to the training data/process, while SnatchML  operated fully at inference time. Song et al. introduced the notion of overlearning [42], proposing its use to perform repurposing attacks. Their method involves training a separate adversary model to infer sensitive attributes from a trained model’s representations. This adversary model is trained on labeled data to predict sensitive attributes based on extracted representations. Additionally, [42] demonstrates how a trained model can be reused to predict a different sensitive attribute by fine-tuning a new classifier added to an existing model layer. In contrast to SnatchML , [42] requires additional training and labeled data, making it resource-intensive. Moreover, their approach operates in a white-box setting, whereas SnatchML can function effectively in a black-box scenario.
Transfer Learning and Adversarial Reprogramming: Transfer learning (TL) approaches involve fine-tuning a model fully or partially to adapt to a new domain, leveraging the knowledge gained from the source domain [47]. SnatchML differs from standard TL in that it does not require model fine-tuning. This is especially beneficial in black-box scenarios, such as cloud infrastructures, where attackers do not have access to gradients or internal model components. SnatchML operates in cases where transfer learning is not possible, allowing attackers to abuse model outputs without directly interacting with the model.
AR operates at test time; repurposes a pretrained neural network for a new task by applying a specific transformation to the input data, effectively mapping it from the target domain to the source domain [13]. This transformation, or adversarial program, is optimized so that the network’s output for the transformed inputs aligns with the desired outputs of the new task. Specifically, the input transformation and output mapping are designed to replace the transfer learning, effectively hijacking the model’s capabilities by manipulating the input and output but without altering its parameters. Some AR approaches include training a neural network for the output mapping. For example, in [26, 34] the approach consists of adding a trainable dense layer between the source model’s output and the target model’s output, which shifts AR closer to a transfer-learning paradigm.
Our approach has several key distinctions with AR: First, similar to other hijacking attacks, AR is constrained by the number of classes present in the victim model, as it repurposes the model’s existing outputs for the new task, unless a TL approach is used. In contrast, SnatchMLhandles a higher number of classes for the hijacking task, offering greater flexibility. Second, AR typically requires white-box access to the model to leverage gradients for optimization. This limits its applicability in settings where only black-box access is available. SnatchML, however, can be executed effectively in both white-box and black-box settings. Finally, AR requires modifying the input, which can be impractical when the attacker, who is also the model owner, aims to covertly misuse a benign model for malicious purposes. In such scenarios, altering the model or its inputs could raise suspicion, rendering AR unsuitable. In contrast, our approach operates without any modifications to the input or the model, enabling more discreet exploitation. We include a comparative analysis with AR in Appendix A-A.
XIV Conclusion
This paper introduces SnatchML, a novel framework for hijacking ML models at inference time, demonstrating a novel threat model for model hijacking. By exploiting the output of the victim model during inference, SnatchML effectively repurposes models for unintended tasks without requiring access to training data and without reprogramming in the input space. Our experiments reveal that SnatchML can perform hijacking tasks even when they are unrelated to the original model’s purpose, highlighting the increased risk and potential for harm with limited access privileges. Additionally, we investigate the role of over-parameterization in facilitating these attacks.
Appendix A Appendix
A-A Comparison with Adversarial Reprogramming
Figure 15 shows the results of SnatchML  comparatively with adversarial reprogramming. To be able to establish a comparison with adversarial reprogramming, we only provide results on the feasible combinations of original-hijacking tasks, i.e., where the number of hijacking task classes is less than the number of original task classes.
Figure 15: Comparison between the hijacking attack performance with SnatchML, Adversarial Reprogramming [13], and Unconstrained upper bound [38].
Figure 16: Correlation distribution between output feature maps from pretraine Emotion Recognition and Face Recognition.
A-B Features Correlation
We posit that the models might learn more than they should in the training process, including learning other tasks. Our intuition is that ML models, perhaps due to over-parametrization, have a by-design capacity that exceeds the minimum necessary capacity to learn the task they are trained for. While it has been shown in the literature that models can overfit and memorise data making them vulnerable to certain attacks such as Membership [48] and property inference [16] attacks, the risk that ML models (unintentionally) generalize to other tasks is yet to be explored/exploited. In this section, we provide a preliminary analysis to gain insights in the plausibility of this hypothesis. Given a model $f^{(i)}(.)$ trained on a data distribution $(X,Y)\sim\mathcal{D}$, using a loss function $\mathcal{L}_{i}$, relevant to a Task $\mathcal{T}_{i}$. Let $f^{(j)}(.)$ be a model with the same architecture as $f^{(i)}(.)$, but is trained on a data distribution $(X’,Y’)\sim\mathcal{D}’$ ’, using a loss function $\mathcal{L}_{j}$, relevant to a Task $\mathcal{T}_{j}\neq\mathcal{T}_{i}$. We ask whether features learnt by $f^{(i)}(.)$ are correlated with those learnt by $f^{(j)}(.)$? We train two ResNet-9 and two Mobilenet architectures each for both Emotion Recognition and Face Recognition. We then check the correlation coefficient distribution between the same layers’ features maps that correspond to the 2 tasks. The correlation coefficient between two variables $X$ and $Y$, denoted as $r$, is defined as:
$r=\frac{\sum_{i=1}^{n}(X_{i}-\overline{X})(Y_{i}-\overline{Y})}{\sqrt{\sum_{i=1}^{n}(X_{i}-\overline{X})^{2}}\sqrt{\sum_{i=1}^{n}(Y_{i}-\overline{Y})^{2}}}$ (6)
where:
  • $n$ is the number of data points,
  • $X_{i}$ and $Y_{i}$ are the individual values of the variables $X$ and $Y$,
  • $\overline{X}$ and $\overline{Y}$ are the means of $X$ and $Y$ respectively.
Figure 16 shows the features correlation distribution for Layer 5 and layer 11 for ResNet-9, and Layer 2 and Layer 4 for Mobilenet. Intrestingly, Figure 16 shows a positive correlation suggesting a potential overlap between what the model learns for tasks that share semantic clues. In other words, this preliminary observation supports the idea that models trained on an original task are capable of extracting features relevant to other tasks.
A-C Impact of limited output classes
In continuation of our study on the performance of the SnatchML attack, we now explore how varying the number of top logits impacts the effectiveness of repurposing a pre-trained model for a new classification task. This analysis investigates the use of different subsets of the highest-ranked logit values from a pre-trained model originally intended for a different task. The motivation behind this study lies in understanding how the selection of top-ranked logit values influences the effectiveness of the SnatchML attack. By ranking and selecting the top k logits from the model’s outputs, we investigate how accurately the model can be hijacked for a new classification task.
We illustrate here three distinct cases where a pretrained MobileNet model, originally trained on ImageNet, is adapted for ethnic classification using the UTKface dataset, pneumonia classification using chest X-ray images and face recognition using Olivetti dataset. Our investigation focuses on how different levels of high-confidence predictions (top-k logits) influence the effectiveness of the SnatchML attack. By varying the parameter k (representing the number of top-ranked logits used), we explore the model’s vulnerability to repurposing for these new tasks, which were not part of its original training scope. The experimental results are presented across different $k$ values—specifically, $k=1$,$10$ ,$100$ , and $1000$. For each $k$, the accuracy metrics reveal the model’s performance in correctly predicting ethnicity and pneumonia.
Across the three datasets, the results, presented in Figure 17 exhibit a consistent trend: as k increases, the model’s accuracy improves. This trend indicates that the critical information required for the SnatchML attack—utilizing the model’s ability to classify ethnicities, pneumonia and faces is concentrated within the top logits of its prediction outputs. Notably, there is a noticeable convergence in accuracy metrics as k increases, particularly evident from k=100 onwards. This convergence suggests that a significant portion of the useful information for our attack is captured within the top $10\%$ of the highest-ranked logits. This finding underscores that a relatively small subset of confident predictions suffices to effectively repurpose the model for these unintended classification tasks. Furthermore, additional experiments confirming similar observations are presented in the appendix.
Figure 17: ”Effectiveness of SnatchML Attack: Impact of Top Logits Selection Across Diverse Hijacking Tasks.
A-D Impact of the hijacking task’s complexity
To learn more about how task complexity affects attack performance, we ran an additional experiment in both the Olivetti and CK+ datasets, varying the number of identities. This enabled us to analyze how the SnatchML assault behaved as the secondary task’s complexity rises, offering more insight into the relationship between task difficulty and attack efficacy. The Figure 18 represents the relationship between Top-1 accuracy and the number of identities for the two datasets, highlighting the impact of dataset complexity on model performance in a controlled threat model scenario with a ResNet-18 model trained for ER. Top-1 accuracy for both datasets drops as the number of identities increases, which is to be expected given that the task gets fundamentally more difficult with more identities. The Olivetti dataset consistently exhibits higher Top-1 accuracy than CK+, indicating that the identities in Olivetti are easier for the model to distinguish. This difference can be attributed to the greater similarity between classes in CK+, where subtle variations make accurate predictions more difficult. The intraclass/interclass variation ratios support this observation; Olivetti’s ratio of 0.65 suggests relatively distinct identities, while CK+’s higher ratio of 0.88 indicates that intraclass variation is closer to interclass variation, reflecting increased complexity. Within the black-box, inference-only threat model, where the attacker lacks direct input access, this analysis helps explain why accuracy is particularly low for complex datasets like CK+. The results demonstrate that the subtle differences between classes in datasets with higher complexity create significant challenges for the model, especially when direct access to input manipulation or gradient feedback is unavailable.
Figure 18: Exploring the Relationship Between Complexity (Identity Count) and Attack Efficiency (Top-1 Accuracy.)
A-E Model Hijacking for Unrelated Tasks
To investigate the generalizability of model hijacking to tasks unrelated to the original training task, we conducted experiments using ResNet-18, MobileNet, and ResNet-50. These models, originally trained on ImageNet or CIFAR-10, were evaluated on a variety of unrelated tasks, including MNIST, CIFAR-100, CK+ (emotion and identity recognition), Olivetti, the Celebrity dataset, and a synthetic dataset. Results are shown in Tables VI, VII, VIII and IX.
TABLE VI: Hijacking attack on ResNet-18 Trained on Unrelated Tasks (ImageNet) Across Multiple Applications
Application Random Guess Trained on ImageNet (Fv) Trained on ImageNet (logit)
MNIST 10% 93% 91%
CIFAR100 1% 23% 20%
CK+ emotions 16.6% 56% 60%
CK+ identities 1.1% 65% 63%
Olivetti 2.5% 91% 87%
Celebrity 11% 100% 100%
Synthetic 2.1% 100% 100%
TABLE VII: Hijacking attack on MobileNet Trained on Unrelated Tasks (ImageNet) Across Multiple Applications
Application Random Guess Trained on ImageNet (Fv) Trained on ImageNet (logit)
MNIST 10% 86% 85%
CIFAR100 1% 15% 14%
CK+ emotions 16.6% 48% 53%
CK+ identities 1.1% 68% 57%
Olivetti 2.5% 76% 73%
Celebrity 11% 100% 100%
Synthetic 2.1% 100% 100%
TABLE VIII: Hijacking attack on ResNet-50 Trained on Unrelated Tasks (Cifar-10) Across Multiple Applications
Application Random Guess Trained on CIFAR-10 (Fv) Trained on CIFAR-10 (logit)
MNIST 10% 91% 62%
CIFAR100 1% 13% 6%
CK+ emotions 16.6% 48% 48%
CK+ identities 1.1% 65% 45%
Olivetti 2.5% 94% 64%
Celebrity 11% 100% 100%
Synthetic 2.1% 99% 99%
TABLE IX: Hijacking attack on MobileNet Trained on Unrelated Tasks (Cifar-10) Across Multiple Applications
Application Random Guess Trained on CIFAR-10 (Fv) Trained on CIFAR-10 (logit)
MNIST 10% 58% 57%
CIFAR100 1% 7% 6%
CK+ emotions 16.6% 44% 51%
CK+ identities 1.1% 49% 49%
Olivetti 2.5% 49% 43%
Celebrity 11% 100% 100%
Synthetic 2.1% 99% 99%
To further illustrate our finding, Figure 19 illustrates the t-SNE distribution of identity classes for three datasets (Olivetti, Celebrity, and Synthetic) based on their feature vectors derived from inference using a ResNet18 model pretrained on ImageNet. The figures distinctly exhibit the existence of separate clusters corresponding to each identity. This finding further reinforces the notion of model hijacking’s potential, as it demonstrates that pretrained models can capture meaningful representations even for tasks unrelated to their original training, echoing the observations made in the preceding paragraph.
Figure 19: t-SNE distributions for identity classes across three datasets (Olivetti, Celebrity, and Synthetic) using ResNet18 pretrained on ImageNet, based on their feature vectors. The Euclidean distance illustrates the closest data points: solid lines connect data points of the same identity, while dashed lines connect data points of different identities.
Table X presents the results of additional experiments evaluating the performance of randomly initialized neural networks across various applications, further illustrating the capabilities of models in hijacking tasks that are entirely unrelated to their original training.
TABLE X: Hijacking attack on a randomly initialized 2D-CNN model to perform ER, Identification on CK+ and image classification on MNIST and CIFAR100
Application Random Guess Trained on CK+ (Fv) Trained on CK+ (logit)
Emotion (CK+) 16.6% 48% 42%
Identity (CK+) 1.1% 67% 47%
Cifar100 1% 5% 2.3%
MNIST 10% 80% 39%
A-F Assessing Model Hijacking Across Additional Qualitative Results
In this section, we present additional qualitative results for the scenario discussed in Section V. In Figure 20, we depict some examples of predictions made by an Ethnic Groups trained model used for predicting age groups on the UTKface dataset. The five nearest images are displayed for each query. For instance, in the first row, despite the fact that the two closest images do not belong to the correct age group (old), they are both individuals within the age group boundary (adult). Additionally, we observe in the third row that both errors belong to the nearest age group, which could be considered as containing relatively useful information. In Figure 21, we present some examples of results obtained from a Gender Groups trained model used for predicting ethnic groups on the UTKface dataset. The five nearest images are displayed for each query. In the second row, we observe that the model returns images with similar attributes such as mustaches or glasses, which are unrelated to the original task it was trained for, namely predicting Gender groups. This result illustrates the fact that the model unintentionally learns additional information that can be exploited in unauthorized ways.
Figure 20: Ethnic groups trained model used to predict age group on UTKface dataset. The 5 nearest images are displayed for each query.
Figure 21: Gender groups trained model used to predict ethnic group on UTKface dataset. The 5 nearest images are displayed for each query.